Massachusetts Institute of Technology Handout 1 Mathematics for Computer Science August 1, 2005 Professor Charles E. Leiserson

Minicourse Information

MIT Subject

This two-lecture minicourse is based on the MIT subject 6.042J/18.062J Mathematics for Computer Science, which covers discrete mathematics and probability. The material in the MIT subject 6.042 is a prerequisite for SMA5503. In Spring 2003, videos and course materials for 6.042 were compiled as SMA5508 and are available from the SMA website.

Staff

MIT Instructor Charles E. LEISERSON cel@mit.edu

Schedule

Lectures are held in room 1-390. Please plan to arrive on time.

Proofs, Induction Monday, August 1 1:00 – 2:30
Probability  Tuesday, August 2 1:00 – 2:30

How to Learn More

Two textbooks are used for the full-length version of this course at MIT. The book The Nuts and Bolts of Proofs by Antonella Cupillari is a small paperback with content similar to our first lecture. The book Discrete Mathematics and Its Applications by Kenneth Rosen is an 800-page text. Sections 3.1 and 3.2 of Rosen’s book cover our first lecture, and sections 4.4 and 4.5 (very briefly) cover the second and third. You must know all the material in 6.042 as a prerequisite for SMA5503, even though we will be reviewing only part of the material.